The pitch P of a musical tone varies inversely as its wavelength W. One tone has a pitch of 283 vibrations per second and a wavelength of 12 ft. Fine the wavelength of another tone that has a pitch of 348 vibrations per second.
Answer by Alan3354(69443)

The pitch P of a musical tone varies inversely as its wavelength W. One tone has a pitch of 283 vibrations per second and a wavelength of 12 ft. Fine the wavelength of another tone that has a pitch of 348 vibrations per second.
--------------
p*w = k (some constant)
P*w = 283*12 = 3396
----------
w = 3396/p
w = 3396/348
w = 9.7586 feet
------------------
PS I would call it frequency, not pitch